### v2

# arXiv:2511.10155v2[quant-ph] 02 Mar 2026

Abstract
Quantum backflow is a counterintuitive phenomenon in which the probability density of a quantum particle propagates opposite to its momentum. Experimental observation of backflow has remained elusive due to two main challenges: (i) the effect is intrinsically small, with less than 4% of the probability able to flow backward, and (ii) it requires wave packets with a well-defined momentum direction, which are difficult both to prepare and to verify under realistic, noisy conditions. Here, we overcome these challenges by introducing a general formulation of quantum backflow applicable to arbitrary momentum distributions. The framework recovers the standard backflow limit for unidirectional states and identifies general backflow as probability flow exceeding that predicted by the particle’s momentum distribution alone. We show that this excess can reach nearly 13%, surpassing the standard backflow bound by more than a factor of three. Furthermore, we extend the framework to the closely related phenomenon of quantum reentry, provide explicit examples of quantum states exhibiting large general backflow and reentry, and discuss the foundational implications of these nonclassical effects. Our results open a pathway toward the experimental observation of quantum backflow in realistic settings.
Introduction. Interference of matter waves—evidenced by the appearance of characteristic density fringes obtained through repeated position measurements of a particle at a given time—has been a cornerstone of quantum mechanics since its early days. Only relatively recently, however, has it been recognized that performing position measurements at multiple distinct times can reveal nonclassical aspects of probability transport even for a free nonrelativistic particle. Perhaps the most well-known example is quantum backflow [1, 2, 3, 4, 5], but other notable effects include quantum reentry [6], Tsirelson’s precession [7, 8, 9, 10, 11, 12], and quantum advantage in various transportation tasks [13, 14, 15, 16]. These effects appeal to the very core of quantum mechanics, much like interference and diffraction. Yet even quantum backflow, which manifests as a classically impossible transfer of probability against the particle’s momentum, has not yet been experimentally observed. [^1] There has been significant experimental progress in studying optical analogs of quantum backflow, where the instantaneous local momentum of a light field is shown to point opposite to the local momentum of its constituent modes. Such effects have been demonstrated in linear motion using classical optical beams [17, 18], and in rotational motion using both classical light [19] and single photons [20]. However, because these experiments are effectively time-independent, they do not capture the probability dynamics that fundamentally define quantum backflow.
Two main obstacles have hindered the experimental observation of quantum backflow. The first arises from the intrinsic weakness of the effect. Bracken and Melloy [4] established a limit to the probability that can flow against the particle’s momentum—the Bracken–Melloy constant $c_{\text{BM}}$—which they estimated to be about $4\%$. This estimate has been progressively refined [21, 22], with the most accurate value being $c_{\text{BM}}=0.0384506$ [23]. (Exact but not optimal analytical bounds on $c_{\text{BM}}$ are reported in Refs. [16, 10].) However, the state saturating this limit is highly singular: it possesses a discontinuous momentum distribution, infinite energy, and diverging probability current, making it physically unrealistic. More regular wave functions [4, 21, 24, 25, 26, 27, 28, 29] yield substantially smaller backflow values, typically below $0.016$, rendering the effect challenging to detect.
The second obstacle is the strict requirement of a fixed momentum direction. Such states are difficult to engineer and verify because realistic, noisy measurements inevitably produce spurious outcomes that can indicate an incorrect momentum direction even for properly prepared initial states. Moreover, in some systems—such as ballistic electrons in mesoscopic conductors—preparing a strictly unidirectional wave packet is fundamentally impossible [30]. While Bracken has generalized the standard backflow formulation to systems with any momentum cutoff [5], these considerations motivate the development of a broader framework capable of describing quantum backflow for arbitrary momentum distributions.
Here, we develop such a general formulation, applicable to both quantum backflow and the closely related phenomenon of quantum reentry. The latter describes the nonzero probability that a free quantum particle reenters a spatial region it has previously left—a behavior forbidden to classical free particles, whose motion is unidirectional in the absence of external forces. Backflow emerges as a special case of reentry [31]. We show that, when extended to arbitrary momentum distributions, quantum reentry can be tested by examining only a finite region of the experimental interferogram, making this approach particularly suitable for experimental implementation.
Our approach is founded on the observation that quantum interference enables a particle to reach spatial regions inaccessible to classical particles. We express the classical limits on probability transport as two inequalities—one for backflow and one for reentry—that make no assumptions about momentum direction or spatial localization. Quantum mechanics can violate these inequalities, with up to $0.1281$ of the particle’s probability flowing opposite to the classical expectation. The optimal state yielding this maximal violation differs from the standard backflow-maximizing state. We also investigate general backflow in experimentally relevant settings using superpositions of Gaussian wave packets. While Gaussian states have been considered in prior studies of standard backflow [24], only the present work provides a rigorous treatment that accounts for their infinite momentum support.
Classical limits. We begin by characterizing the absence of backflow in classical systems. Consider a free classical particle of mass $m$ moving along the $x$ axis, with an unrestricted phase-space distribution f(x,p,t)=f(xpt/m,p,0)f(x,p,t)=f(x-pt/m,p,0). The central quantity in backflow-like problems is the probability of finding the particle to the left of the origin $x=0$ at time $t$, denoted as $P_{-}(t)=\operatorname{Prob}(x<0,t)$. This probability can change between times $t_{1}$ and $t_{2}=t_{1}+T$ in two distinct ways: (i) particles with positive momentum ($p$ > 0) initially located in the interval $x\in(-pT/m,0)$ will leave the region, while (ii) particles with negative momentum ($p<0$) initially positioned in $x\in(0,-pT/m)$ will enter it. Since in general backflow we are interested in the increment of this probability, we upper-bound the probability gain by the second contribution only:
$\hskip-8.5359ptP_{-}(t_{2})-P_{-}(t_{1})\leq\int_{-\infty}^{0}\!\!dp\int_{0}^{-\frac{pT}{m}}\!\!dx\,f(x,p,t_{1})\leq\tilde{P}_{-}\,$. (1)
Here, $\tilde{P}_{-}=\operatorname{Prob}(p<0)$ is the probability of negative momentum, which remains time independent under free evolution. The second inequality follows by extending the $x$-integration limits to the entire real line and using the non-negativity of the probability density $f$.
Inequality (1) is expressed solely in terms of measurable position and momentum marginals, making it directly comparable to quantum results. For $\tilde{P}_{-}=0$, it reduces to the familiar statement that the probability of finding a right-moving particle to the left of the origin cannot increase with time. In its general form, however, the inequality makes no assumptions about the marginals and simply bounds the position-probability gain by the probability of negative momentum. Any violation of this bound—shown below to be possible for quantum wave packets—therefore signals a nonclassical transport effect where the particle moves backward more than is permitted by its momentum distribution.
The momentum probability in (1) can be traded for a position probability to produce a generalization of the reentry problem [6]. Consider a finite interval along the $x$ axis, r=(s,0)r=(-s,0) for concreteness, and define $P_{r}(t)=\operatorname{Prob}(x\in r,t)$ as the probability for the particle to be inside $r$ at time $t$. The difference $P_{r}(t_{2})-P_{r}(t_{1})$ quantifies the fraction of particles that have entered $r$ between times $t_{1}$ and $t_{2}>t_{1}$. Since classical particles can only enter from outside the region, this increase cannot exceed the total fraction of particles that were outside $r$ at any earlier time $t_{0}<t_{1}$, namely $1-P_{r}(t_{0})$. This leads to the following inequality characterizing the impossibility of reentry for classical particles:
$P_{r}(t_{2})-P_{r}(t_{1})\leq 1-P_{r}(t_{0})\,$, (2)
which holds for any $t_{0}<t_{1}<t_{2}$. [^2] See Supplemental Material for an alternative derivation of Eqs. (1) and (2). For $P_{r}(t_{0})=1$ and $s\to\infty$, it reduces to the classical reentry condition discussed in Refs. [6, 31].
Compared to (1) the reentry inequality offers two notable advantages. First, it depends solely on position measurements, allowing it to be tested experimentally by comparing interference patterns at three distinct times. Second, only a finite region of the interferogram, corresponding to $r$, needs to be analyzed, provided the total flux of particles is known.
Simple example. We now demonstrate that the classical bounds expressed in inequalities (1) and (2) can be violated by quantum particles. Consider a wave function consisting of a superposition of two Gaussian wave packets, $\Psi(x,t)=\sum_{n=1,2}c_{n}g_{n}(x,t)$, directly accessible in optical systems and Bose-Einstein condensates (BECs). This type of superposition was previously analyzed in Ref. [24] for real coefficients $c_{n}$ and Gaussians $g_{n}(x,t)$ with identical widths, yielding quantum backflow of $0.0061$. Although strictly nonzero, the corresponding probability of negative momentum was found to be small, $\tilde{P}_{-}\approx 10^{-10}$, and was therefore neglected.
To systematically study the effect of negative momentum, we define
$\Delta_{\mathrm{QB}}$ $=$ $P_{-}(t_{2})-P_{-}(t_{1})-\tilde{P}_{-}\,$, (3)
$\Delta_{\mathrm{RE}}$ $=$ $P_{-}(t_{2})-P_{-}(t_{1})+P_{-}(t_{0})-1\,$, (4)
which, according to Eqs. (1) and (2), cannot exceed zero in classical mechanics. Details of maximizing these quantities over the set of superposed Gaussian states are provided in the Supplemental Material (SM). Using the same assumptions as in Ref. [24], we find that the generalized quantum backflow reaches $\Delta_{\mathrm{QB}}=0.0106$. Relaxing the assumptions on the Gaussian superposition increases the maximum backflow to $\Delta_{\mathrm{QB}}=0.0120$. The same enhancements are observed for quantum reentry, $\Delta_{\mathrm{RE}}$. While these values remain modest, they clearly demonstrate that incorporating negative-momentum components allows substantially larger violations of classical probability bounds.
Large backflow example. The previous illustration showed that general quantum backflow can exceed the magnitude observed in the standard scenario. Remarkably, general backflow can even significantly surpass the Bracken–Melloy constant, $c_{\text{BM}}=0.0384506$. Inspired by Ref. [25], we construct an explicit example using the momentum-space wave function $\tilde{\psi}(p)=\Theta(p)a_{1}(b_{1}-p)e^{-c_{1}^{2}p^{2}}+\Theta(-p)a_{2}(b_{2}-p)e^{-c_{2}^{2}p^{2}}$, where $a$’s, $b$’s, and $c$’s are constants, and $\Theta(\cdot)$ is the Heaviside step function. Such superpositions are feasible in BEC platforms. With a suitable choice of parameters, this state achieves a general backflow of $\Delta_{\mathrm{QB}}=0.0624$. There also exists a corresponding state exhibiting the same value of $\Delta_{\mathrm{RE}}$. Detailed derivations and parameter choices are provided in the SM.
Unified approach. We now demonstrate that the expressions for $\Delta_{\mathrm{QB}}$ and $\Delta_{\mathrm{RE}}$ with the range $r=(-\infty,0)$ can be cast in the same mathematical form. For general backflow, we write the time-dependent wave function of a free particle as
$\Psi(x,t)=\frac{1}{\sqrt{2\pi\hbar}}\int_{-\infty}^{+\infty}dp\,\tilde{\psi}(p)\exp\left(-\frac{it}{2m\hbar}p^{2}+\frac{ix}{\hbar}p\right),$ (5)
where $\tilde{\psi}(p)$ is the normalized momentum-space wave function at $t=0$. The relevant probabilities in Eq. (3) are then
$P_{-}(t)=\int_{-\infty}^{0}dx\,|\Psi(x,t)|^{2}\,,\quad\tilde{P}_{-}=\int_{-\infty}^{0}dp\,\big|\tilde{\psi}(p)\big|^{2}\,.$ (6)
Following the approach of Ref. [4], detailed for convenience in the SM, we find that $\Delta_{\mathrm{QB}}=\Delta$, where
$\Delta$ $=$ $\int_{-\infty}^{+\infty}du\int_{-\infty}^{+\infty}du^{\prime}\,\varphi^{*}(u)K(u,u^{\prime})\varphi(u^{\prime})\,$, (7)
$K(u,u^{\prime})$ $=$ $-\frac{\sin\left(u^{2}-{u^{\prime}}^{2}\right)}{\pi(u-u^{\prime})}-\Theta(-u)\delta(u-u^{\prime})\,$, (8)
and
$\varphi(u)=\left(\frac{4\hbar m}{t_{2}-t_{1}}\right)^{\!\!1/4}\exp\!\left(-i\frac{t_{2}+t_{1}}{t_{2}-t_{1}}u^{2}\right)\tilde{\psi}\!\left(u\sqrt{\frac{4\hbar m}{t_{2}-t_{1}}}\right)\,.$ (9)
Here, $\delta(\cdot)$ denotes the Dirac delta function, and $\varphi(u)$ is a rescaled momentum-space wave function normalized as $\int_{-\infty}^{+\infty}du\,|\varphi(u)|^{2}=1$.
For quantum reentry with $r=(-\infty,0)$, we express $\Psi(x,t)$ via the free-particle propagator
$\Psi(x,t)=\sqrt{\frac{m}{2\pi i\hbar\tau}}\int_{-\infty}^{+\infty}dx^{\prime}\,\exp\!\left(i\frac{m(x-x^{\prime})^{2}}{2\hbar\tau}\right)\psi(x^{\prime})\,$, (10)
where $\tau=t-t_{0}$, and $\psi(x)$ is the position-space wave function at $t=t_{0}$. Following the method of Ref. [6], detailed in the SM, $\Delta_{\mathrm{RE}}$ can be expressed in the same form, $\Delta_{\mathrm{RE}}=\Delta$, with $\varphi(u)$ now representing a rescaled position-space wave function:
$\varphi(u)$ $=$ $\left(\frac{4\hbar\tau_{2}\tau_{1}}{m(\tau_{2}-\tau_{1})}\right)^{\!\!1/4}\exp\!\left(i\frac{\tau_{2}+\tau_{1}}{\tau_{2}-\tau_{1}}u^{2}\right)$ (11)
$\times$ $\psi\!\left(-u\sqrt{\frac{4\hbar\tau_{2}\tau_{1}}{m(\tau_{2}-\tau_{1})}}\right)\,$,
where $\tau_{1}=t_{1}-t_{0}$ and $\tau_{2}=t_{2}-t_{0}$.
Thus, $\Delta$ captures the probability transfer in both the general backflow and reentry scenarios. Compared with the standard backflow [4] and reentry [6] formulations, two key differences appear: the integrals over $u$ and $u^{\prime}$ extend over the entire real line, and the kernel $K(u,u^{\prime})$ includes an additional term, $\Theta(-u)\delta(u-u^{\prime})$. These modifications account for the larger values of $\Delta$ observed in the general setting. In the SM we also give a perturbative argument, based on modifications of the state maximizing standard positive-momentum backflow, showing that $\Delta$ can exceed the Bracken–Melloy limit.
Supremum of $\Delta$. The maximal quantum violation of the classical inequalities (1) and (2) is obtained by optimizing $\Delta$ over all normalized functions $\varphi(u)$. This is achieved by performing an unconstrained optimization of $\Delta+\lambda\int_{-\infty}^{+\infty}du\,|\varphi(u)|^{2}$, where $\lambda$ is a Lagrange multiplier [4]. The resulting Euler-Lagrange equation reads
$\int_{-\infty}^{+\infty}du^{\prime}\,K(u,u^{\prime})\varphi(u^{\prime})=\lambda\,\varphi(u).$ (12)
so that the supremum of $\Delta$ corresponds to the largest eigenvalue of the integral operator: $\sup_{\varphi}\Delta=\sup\{\lambda\}$.
To evaluate $\sup_{\varphi}\Delta$ numerically, we first truncate the integration domain in Eq. (12) to (L,L)(-L,L), giving eigenvalues $\lambda_{L}$ such that $\lambda=\lim_{L\to\infty}\lambda_{L}$. We then discretize (L,L)(-L,L) into $2N$ intervals, replacing the integral with a finite sum, thereby converting the integral eigenproblem into a matrix eigenproblem with eigenvalues $\lambda_{L,N}$, which satisfy $\lim_{N\to\infty}\lambda_{L,N}=\lambda_{L}$.
In our calculations, $L$ was varied from $10$ to $40$ in steps of $5$, and for each $L$, NN was increased from hundreds to thousands to determine the maximal $\lambda_{L,N}$. For fixed $L$, the maximal $\lambda_{L,N}$ scales approximately linearly with $1/N$; we therefore estimate the limit $N\to\infty$ and obtain $\lambda_{L}$ by fitting a linear function to the computed eigenvalues. The optimal fit parameters and associated uncertainties are obtained via standard linear regression [32]. Similarly, the maximal $\lambda_{L}$ scales approximately linearly with $1/L$, allowing us to estimate the limit $L\to\infty$ and its uncertainty. The resulting value, with one standard deviation error, is
$\sup\Delta=0.128100\pm 0.000002\,.$ (13)
Comprehensive details of the numerical analysis are provided in the SM.
Figure 1: Rescaled state maximizing general backflow and reentry compared with the standard (positive-momentum) backflow-maximizing state.
The eigenvector corresponding to the largest eigenvalue provides an approximation of the state $\varphi(u)$ that maximizes general backflow or reentry. As shown in Fig. 1 this state (solid orange curve) exhibits a discontinuity at $u=0$ and long oscillatory tails extending to infinity, similar to the standard backflow-maximizing state (dashed blue curve for positive-momentum problem).
[…]
The best known analytical bounds on the maximal eigenvalue of $\Omega$ were established by connecting it to Tsirelson’s precession problem [7], and applying very recent results of Ref. [8, 10]. This shows that $0.128092\leq\sup\Delta\leq 0.192466$, exceeding the Bracken-Melloy bound $c_{\text{BM}}\simeq 0.0384506$ by more than a factor of three, and with the lower bound coming remarkably close to our estimate, given by Eq. (13).
Discussion. We conclude with remarks on the foundational implications of our results. In the general backflow formulation, we considered the probability of locating the particle to the left of the origin. Naturally, the origin can be replaced by any point $b$ on the $x$ axis. Quantum backflow occurs whenever $\Delta_{\text{QB}}=\operatorname{Prob}(x<b,t_{2})-\operatorname{Prob}(x<b,t_{1})-\tilde{P}_{-}$ [c.f. Eq. (3)] becomes positive. Similarly, one can consider quantum “overflow”, recently introduced in Ref. [23], which occurs when $\Delta_{\text{QO}}=\operatorname{Prob}(x>f,t_{2})-\operatorname{Prob}(x>f,t_{1})-\tilde{P}_{+}$ becomes positive, with $f$ an arbitrary point and $\tilde{P}_{+}=1-\tilde{P}_{-}$ the probability of positive momentum. Since in the general formulation introduced here the particle can move in any direction, a natural question arises: can a particle simultaneously exhibit backflow and overflow? In other words, does there exist a wave function for which $\Delta_{\text{QB}}>0$ and $\Delta_{\text{QO}}>0$ for some choice of parameters? The answer is negative. It can be shown (see SM) that $\Delta_{\mathrm{QB}}+\Delta_{\mathrm{QO}}\leq 0$ for all parameters, revealing a trade-off and mutual incompatibility between these two effects. It is also worth noting that while quantum overflow can occur only over multiple disjoint time intervals in states with strictly positive momentum [23], it can already arise within a single interval for states with an arbitrary momentum distributions. In fact, quantum backflow and overflow are qualitatively equivalent phenomena, distinguished only by the direction of the probability transport.
The Bracken–Melloy constant $c_{\mathrm{BM}}$ characterizes maximal standard (positive-momentum) backflow and is independent of the particle’s mass, the backflow duration, and Planck’s constant. It was proposed to reflect the mathematical structure of quantum free-particle evolution [4]. The same properties hold for the general backflow constant $\sup\Delta$, given by Eq. (13), suggesting that quantum free evolution is characterized by at least two fundamental dimensionless constants. Understanding the relation between them presents a promising direction for future research.
A more pragmatic follow-up question is how the strength of general backflow changes when the experimental evaluation of $\operatorname{Prob}(x<0,t)$ uses a smoothed spatial window rather than a sharp one. In the case of the standard positive-momentum formulation, such smoothing is known to lead to an underestimation of quantum backflow [24]. Another important aspect of a future experimental validation of general backflow is ensuring that any violation of inequalities (1) or (2) arises from quantum interference rather than from a failure of the dynamical assumptions (e.g., force-free motion) of the system under investigation. A theoretical framework for eliminating such a dynamical loophole in the context of the Tsirelson precession problem has recently been developed in Ref. [12].
The classical limits, Eqs. (1) and (2), rely on two assumptions: (i) the existence of a joint probability density $f(x,p,t)$, and (ii) the validity of classical equations of motion for a free particle, $x_{t}=x_{0}+pt/m$ and $p_{t}=\text{const}$. Quantum backflow, reentry, and other nonclassical probability transport effects violate at least one of these assumptions. Relaxing (i) while keeping (ii) leads to pseudo-probability frameworks, such as the Wigner function formalism, in which backflow emerges as negative probability transfer of particles in straight line motion [4, 24]. Conversely, retaining (i) and relaxing (ii) leads to Bohmian mechanics, where even free particles follow curved trajectories due to an effective nonlocal potential [33, 22]. Yet, these trajectories have to comply with the Ehrenfest theorem. Moreover, one can demonstrate that any covariance measurements remain compatible with the straight line motion. Consequently, any signatures of backflow can only appear in the dynamics of third and higher moments of the position distribution. This demonstrates both the subtlety of the effect and the sensitivity of our methods to these higher-order changes.
We also note that the derivation of the inequalities (1) and (2) does not require strictly free-particle equations of motion. Rather, they hold for any continuous phase-space flow satisfying $\operatorname{sgn}(dx_{t}/dt)=\operatorname{sgn}(p_{t})$ and $\operatorname{sgn}(p_{t})=\text{const}$, where $\operatorname{sgn}(\cdot)$ denotes the sign function. These conditions require that the particle’s velocity be aligned with its momentum and that the direction of motion be preserved. This observation opens the possibility of investigating general quantum backflow and reentry beyond the free-particle setting.
In summary, the methods developed here provide a realistic route toward experimental observation of quantum backflow and reentry, while raising several fundamental and practical questions for further exploration.
Acknowledgments. The authors thank Anat Daniel, Bohnishikha Ghosh, Bernard Gorzkowski, Thomas Juffmann, Radosław Łapkiewicz, Timon Zipfelmaier for insightful discussions.